OpenStudy (anonymous):

6h^2+17h+12=0 Solve Solving Equations by Factoring

OpenStudy (anonymous):

6h^2+17h+12=0 (2x+3)(3x+4)=0 (2x+3)=0 2x=-3 x=(-3/2) (3x+4)=0 3x=-4 x=(-4/3)